# Interplay between topology and interactions in superconducting chains

### v1

@@ Abstract
Most studies of non-trivial topological systems are carried out in non-interacting models that admit an exact solution. This raises the question, to which extent the consideration of electronic correlations, present in real systems, modify these results. Exact solutions of correlated electronic systems with non-trivial topological properties, although fundamental are scarce. Among the non-interacting soluble models, we single out the Kitaev p-wave superconducting chain. It plays a crucial role in clarifying the appearance of emergent quasi-particles, the Majorana modes, associated with non- trivial topological properties. Given the relevance of this model, it would be extremely useful if it can be extended to include correlations and still remain solvable. In this work we investigate a superconducting Kitaev chain that interacts through a Falicov-Kimball Hamiltonian with a background of localized electrons. For some relevant values of the parameters of this extended model solutions can be obtained, which allows for a detailed study of the interplay between electronic correlations and non- trivial topological behavior. Besides, the random occupation of the chain by the local moments brings the new interesting effects associated with disorder.
@@ I Introduction
Systems with non-trivial topological properties have raised enormous interest in recent years [1, 2, 3, 4, 5]. They have expanded the landscape of interesting materials in the periodic table, giving rise to new concepts and to the discovery of unsuspected physical properties. Theoretical study of models with non-trivial topological properties has in many cases anticipated unusual behavior creating exciting challenges to experimentalists. One clear example is the Kitaev superconducting chain model, with emergent new quasi-particles, the Majorana fermions in its extremities [6]. The unusual properties of these modes, with possible applications in quantum computers [7, 8, 9, 10, 11, 12] is leading to enormous progress in experimental nano-physics [1, 3, 4, 5, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23].
@@
Many of the models of topological superconductors and insulators are idealized, since they do not include electronic correlations, and consequently can be solved exactly. In one aspect this is very satisfactory, as it allows to obtain a clear picture of the nature of the topological non-trivial properties. On the other hand, it raises the question to which extent these properties are robust to electronic correlations present in real materials [24, 25, 26, 27]. The motivation of the present study is to explore this problem presenting a non-trivial topological interacting model with an exact solution for some values of its parameters. The model we focus in this work is the superconducting $p$-wave chain introduced by Kitaev, extended to include a coupling of the electrons to a background of localized electrons through a Falicov-Kimball (FK) interaction [28, 29]. This many-body problem can be mapped in a non-interacting system with a site dependent chemical potential, which has been extensively studied [31, 32, 32]. Some of these results can be directly used in our analysis, with peculiarities arising from the physical motivation of our model. Our work allows to explore on firm ground the interplay of electronic correlations and non-trivial topology. We consider different occupations of the sites of the chain by localized electrons, including the case of random occupations of these sites that give rise to new and interesting effects.
@@
We organize the paper in the following way: In Sec. II, we introduce the Hamiltonian, which is a generalization of the standard Kitaev chain [6], including a background of localized electrons described by a Falikov-Kimball interaction [28]. In Sec. III, we present the formalism employed in calculating the exact solution of the model. In Sec. IV, we discuss the properties of the standard, spinless, one-dimensional FK model. We consider three cases: Homogeneous or uniform, staggered, and disordered configurations of the localized moments. In Sec. V, we consider the general case, where the conduction band has a p-wave superconducting pairing and interacts with the background of localized $f$-electrons. We present results of the density of states for the homogeneous and staggered cases. Finally, we study numerically the appearance of Majorana zero modes (MZMs) for the finite chain, including a simulation of the disordered case. Finally, in Sec. VI, we give a summary of the results, their physical consequences and new perspectives.
@@ II The Hamiltonian
The Hamiltonian we consider, describes a chain of spinless, itinerant electrons with a $p$-wave pairing that interact with a background of localized electrons through a Falicov-Kimball interaction [28]. The Hamiltonian describing the $p$-wave superconductor is the Kitaev one [1, 6],
$\mathcal{H_{K}}=-\sum_{i,j}t_{ij}c^{\dagger}_{i}c_{j}-\sum_{i,j}\left(\Delta_{ij}c^{\dagger}_{i}c^{\dagger}_{j}+\Delta^{*}_{ij}c_{i}c_{j}\right)-\mu\sum_{i}c^{\dagger}_{i}c_{i},$ (1)
and that of the localized electrons [28],
$\mathcal{H_{L}}=\sum_{i}(E_{f}-\mu)f^{\dagger}_{i}f_{i}.$ (2)
In these equations, $t_{ij}$ is the hopping term of the conduction $c$-electrons, $\Delta_{ij}$ their $p$-wave coupling, $E_{f}$ the energy of the localized $f$-electrons and $\mu$ the chemical potential.
The coupling between the itinerant $c$-electrons and the localized $f$-electrons is of the FK type [28] and given by,
@@
$f^{\dagger}_{i}=(\beta_{Bi}-i\beta_{Ai})/2.$
In terms of these new operators we can write the interaction term as,
@@
As for the other terms in the Hamiltonian, we get,
@@ III The solution
For further progress in the understanding the properties of the model, we consider the case where the chemical potential is at the energy of the localized particles, i.e., $\mu=E_{f}$ and for further simplicity, $E_{f}=\mu$ is taken to be zero. The case $E_{F}=\mu\neq 0$ also can be solved. Notice that for $\mu=0$, the superconducting chain, in the absence of the coupling to the localized electrons, is in a nontrivial topological state, since the condition $|\mu/2t|<1$ is satisfied [1]. On the other hand, for $\Delta=0$ the model corresponds to the one-dimensional spinless FK model [28].
When $E_{f}=\mu=0$, the Hamiltonian $\mathcal{H_{KL}}$ does not depend on the Majorana operators $\beta_{B}$ and $\beta_{A}$, associated with the localized electrons. We can then define an operator $K_{i}=i\beta_{Ai}\beta_{Bi}$ that commutes with the Hamiltonian of the superconductor, i.e., $[K_{i},\mathcal{H_{KL}}]=0$ for all sites. This implies that $K_{i}$ is a $c$-number [26]. On the other hand, the anti-commutation rules of the Majorana operators imply that $K_{i}^{2}=1$ or $K_{i}=\pm 1$. The Hamiltonian of the system can now be written as,
$\mathcal{H}=-\frac{i}{2}\sum_{i}\left[(\Delta+t)\alpha_{Bi}\alpha_{Ai+1}+(\Delta-t)\alpha_{Ai}\alpha_{Bi+1}\right]+i\sum_{i}J_{i}(i\alpha_{Ai}\alpha_{Bi}),$ (8)
where $J_{i}=JK_{i}/16$ ($K_{i}=\pm 1$). From a formal point of view $\mathcal{H}$ is a quadratic form that can be exactly solved. The interacting problem has been mapped in a non-interacting one, formally in a system with a site dependent chemical potential [31, 32, 33].
Let us introduce the quantities
@@
where we wrote $J_{i}=Je^{iQr_{i}}$. The homogeneous, or uniform configuration, with a localized electron on each site of the lattice, $n_{f}=1$, corresponds to $Q=0$. The staggered configuration, where the values of $K_{i}$ alternate between ±$\pm 1$, such that $n_{f}=1$ and $n_{f}=0$, in consecutive sites, corresponds to $Q=\pi$ ($n_{f}=(1/2)(1+i\beta_{Ai}\beta_{Bi})$). Among all possible configurations, these two play an important role, as we discuss below.
The excitations of the system can be obtained from the Hamiltonian of the infinite chain, Eq. 10. For the uniform configuration ($Q=0$), we get,
@@
and $\omega_{h(3,4)}(k)=\pm\omega_{h(1,2)}(k)$. The excitations in the staggered configuration with $Q=\pi$ are given by
@@
and $\omega_{s(3,4)}(k)=\pm\omega_{s(2,1)}(k)$. The condition $\mu=E_{F}=0$ implies that for the non-interacting system, the band of conduction electrons corresponds to a half-filled band of (spinless) conduction electrons. The number of localized $f$-electrons is arbitrary, $n_{f}\leq 1$, and does not affect the total energy of the system in the non-interacting case, $J=0$.
In Fig. 1 we plot the dispersion relations, for specific values of parameters, of the uniform and staggered configurations. For completeness we also show the dispersion of the pure Kitaev chain ($J=0$).
Figure 1: (Color online) Dispersion relations of the excitations in the superconducting state for the a) homogeneous case, $\omega_{h}(k)$, and b) staggered case, $\omega_{s}(k)$, for $t=1$, $\Delta/t=0.2$ and $J/t=0.9$. c) energy of the bogoliubons $\omega_{0}(k)$, of the non-interacting case with the same parameters, but $J=0$ (Kitaev chain).
In Section IV, we will discuss the nature of the superconducting phases in the presence of the interaction with the localized electrons. We investigate how this interaction modifies the topological properties of the Kitaev chain. We performed simulations in finite chains to study the robustness of the Majorana modes in the Kitaev chain to the coupling to the localized fermions. These simulations also yield the ground state energy of different configurations and allow to identify the ground states.
However, before going deeper in the study of the superconducting states, we discuss briefly the spinless FK model in electronic systems. In our model this corresponds to taking the superconducting gap $\Delta=0$ in Eq. 8.
@@ IV The one-dimensional spinless Falicov-Kimball model ($\Delta=0$)
The FK model is well studied in strongly correlated electronic systems, with many applications ranging from metal-insulator to valence transitions [28]. In particular the one-dimensional, spinless FK model has many well-known properties [28]. It can be approached using the technique of bosonisation [28] that yields exact results [28]. From the point of view of topological properties this model is completely trivial.
@@
In this section we briefly discuss the spinless one-dimensional FK model. Formally, we just take the amplitude of the superconducting gap $\Delta=0$ to obtain the dispersion relations for this model. We can distinguish two situations depending on the value of the ratio $(J/t)$.
@@
Figure 2: a) b) c), dispersion relations $\omega_{h}(k)$ of the excitations of the FK model ($t=1$ and $\Delta/t=0$), for the homogeneous case for $J/t=0.75$, $1$ and $1.2$, respectively. d) e) f), dispersion relations $\omega_{s}(k)$ for the staggered case for $J/t=0.75$, $1$ and $1.2$, respectively. In the homogeneous case a) corresponds to a metallic state and at $J/t=1$, (Fig. 2b) there is a metal insulator transition. In the staggered configuration the system is always insulating, unless $J=\Delta$, where it is a Dirac semi-metal. The black line is the energy of the localized quasiparticles that coincides with the chemical potential.
Among the special configurations, we distinguish two cases. For $J/t>1$, both homogeneous and staggered configurations correspond to insulating phases as shown in Fig. 2. Conversely, for $J/t<1$, the homogeneous configuration corresponds to a metallic phase, but the staggered one remains insulating, as can be seen in Fig. 2. The homogeneous configuration corresponds to one localized $f$-electron per site and the staggered to $1/2$ localized electron per site. Since we are working in the grand canonical ensemble, with the condition $E_{f}=\mu=0$, the total energy of the system does not depend directly on the number of $f$-electrons. The dependence on $n_{f}$ is only indirect, since the total energy, given by that of the $c$-electrons is determined by their coupling to the former. In Fig. 3 we show the ground state energies of the special configurations, $Q=0$ and $Q=\pi$, as functions of the interaction $J$. Notice that the staggered configuration has a lower energy compared to the uniform one. In the homogeneous case, with one localized electron per site ($n_{f}=1$), there is a metal-insulator transition at $|J/t|=1$, as shown in Fig. 2 for the dispersion relations.
We point out that the staggered configuration is not the only possibility for a number of localized electrons per site, $n_{f}=1/2$. Unlike the $Q=0$ uniform configuration, there are many other configurations with $<n_{f}>=1/2$, besides the $Q=\pi$, staggered one [29]. These configurations, which are mainly disordered, may become important at high temperatures, as they increase the entropy of the system [28, 29]. In Fig. 3 show the ground state energies, as functions of the interaction $J$, of the special configurations and of random configurations with $<n_{f}>=1/2$. In these disordered configurations there is a probability $p=1/2$ for a site being occupied, or not, by a localized $f$-electron. Notice that the ordered, $Q=\pi$, staggered configuration is always the one with lower energy. Our choice of $\mu=E_{f}=0$ implies that for spinless electrons there is and average number $<N_{c}>=1/2$ of $c$-electrons per site and an arbitrary number of localized $f$-electrons ($n_{f}\leq 1$). The uniform and staggered configurations fixes the number of localized electrons at $n_{f}=1$ and $n_{f}=1/2$, per site, respectively. The former case has a metallic ground state and the latter is an insulator. This closely resembles the Peierls instability where there is a doubling of the unit cell as the system goes from a homogeneous to a staggered configuration [30].
@@
Figure 3: (Color online) Ground state energy of the spinless FK model ($\Delta=0$). Homogeneous case, $n_{f}=1$ (red), staggered, $n_{f}=1/2$ (blue) and random configurations (green) with $<n_{f}>=1/2$. Notice that, the staggered configuration has an energy lower then all the probed random configurations with $<n_{f}>=1/2$.

### v2

@@ Abstract
Most studies of non-trivial topological systems are carried out in non-interacting models that admit an exact solution. This raises the question, to which extent the consideration of electronic correlations and disorder, present in real systems, modify these results. Exact solutions of correlated electronic systems with non-trivial topological properties, although fundamental are scarce. Among the non-interacting soluble models, we single out the Kitaev p-wave superconducting chain. It plays a crucial role in clarifying the appearance of emergent quasi-particles, the Majorana modes, associated with non-trivial topological properties. Given the relevance of this model, it would be extremely useful if it could be extended to include correlations and still remain solvable. In this work we investigate a superconducting Kitaev chain that interacts through a Falicov-Kimball Hamiltonian with a background of localized electrons. For some relevant values of the parameters, this model can be solved exactly by mapping into a non-interacting one. This allows for a detailed study of the interplay between electronic correlations and non- trivial topological behavior. Besides, the random occupation of the chain by the local moments brings new interesting effects associated with disorder.
@@ I Introduction
Systems with non-trivial topological properties have been the subject of intensive studies in recent years [1, 2, 3, 4, 5, 6]. They have expanded the landscape of interesting materials in the periodic table, giving rise to new concepts and to the discovery of unsuspected physical properties. Theoretical study of models with non-trivial topological properties has in many cases anticipated unusual behavior, creating exciting challenges to experimentalists. A clear example is the Kitaev superconducting chain model with emergent new quasi-particles, the Majorana fermions, on its edges [6]. The unusual properties of these modes, with possible applications in quantum computers [7, 8, 9, 10, 11, 12] is leading to enormous progress in experimental nano-physics [1, 3, 4, 5, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23].
@@
Many of the models of topological superconductors and insulators are idealized, since they do not include electronic correlations or disorder, and consequently can be solved exactly. In one aspect this is satisfactory, since it allows to obtain a clear picture of the nature of the emerging quasi-particles and topological non-trivial properties. On the other hand, it raises the question to which extent these properties are robust to electronic correlations, or disorder, present in real materials [24, 25, 26, 27, 28, 29]. The motivation of the present study is to explore this problem presenting a non-trivial topological interacting model with an exact solution for some values of its parameters.
The Falicov-Kimball (FK) model plays an important role in the study of strongly correlated systems, as heavy fermions and mixed-valence material [30, 32, 33]. The model consists of a conduction band of itinerant electrons interacting with a background of localized electrons. In this paper we obtain an exact solution for the one-dimensional spinless FK model with a $p$-wave pairing of the electrons in the conduction band. The solution is possible since we can map the many-body problem in a non-interacting system with a site dependent chemical potential [35, 37, 36, 36]. Our work allows to explore on firm ground the interplay between electronic correlations and non-trivial topology. Since the local moments can randomly occupy the sites in the lattice, we also consider the effect of disorder on the interacting topological system.
@@
We organize the paper in the following way: In Sec. II, we introduce the Hamiltonian, which is a generalization of the $p$-wave superconducting Kitaev chain [6], to include an interaction of the conduction electrons with a background of localized electrons through a Falikov-Kimball Hamiltonian [30]. In Sec. III, we present the formalism and the exact solution of the model. In Sec. IV, we briefly discuss the properties of the pure one-dimensional spinless FK model, since this will be useful to characterize the superconducting phases. Sec. V treats the general case of a conduction band with a $p$-wave pairing that interacts with localized $f$-electrons. We obtain the dispersion relations and density of states of infinite chains. For a full characterization of the topological properties we investigate numerically the appearance of Majorana zero modes (MZMs) in finite chains with open boundary conditions, including the case the local moments are randomly distributed. In Sec. VI, we study in detail random chains, with different probabilities of occupations of the sites by local moments. This allow to examine the effects of disorder in the topological properties. Finally, in Sec. VII we conclude with a summary of our results and a discussion on new perspectives and extensions of the model.
@@ II The Hamiltonian
The Hamiltonian that describes a chain of spinless fermions, with $p$-wave pairing, interacting with a background of localized electrons through a Falicov-Kimball interaction [30], can be written as a sum of three contributions. The first is the Kitaev one [1, 6],
$\mathcal{H_{K}}=-\sum_{i,j}t_{ij}c^{\dagger}_{i}c_{j}-\sum_{i,j}\left(\Delta_{ij}c^{\dagger}_{i}c^{\dagger}_{j}+\Delta^{*}_{ij}c_{i}c_{j}\right)-\mu\sum_{i}c^{\dagger}_{i}c_{i}.$ (1)
The localized electrons [30] are described by,
$\mathcal{H_{L}}=\sum_{i}(E_{f}-\mu)f^{\dagger}_{i}f_{i},$ (2)
where in these equations, $t_{ij}$ is the hopping term of the conduction $c$-electrons, $\Delta_{ij}$ the $p$-wave pairing of electrons in neighboring sites, $E_{f}$ the energy of the localized $f$-electrons and $\mu$ the chemical potential. Finally, the coupling between the itinerant $c$-electrons and the localized $f$-electrons is of the FK type [30] and given by,
@@
$f^{\dagger}_{i}=(\beta_{Bi}-i\beta_{Ai})/2,$
where $\beta_{A,B}^{\dagger}=\beta_{A,B}$. In terms of these new operators we can write the interaction term as,
@@
For the other terms in the Hamiltonian, we get in the Majorana basis,
@@ III The solution
For further progress in understanding the properties of the model, we consider that the chemical potential coincides with the energy of the localized quasi-particles, i.e., $\mu=E_{f}$. We also assume that $\mu=E_{f}=0$, although the case $E_{F}=\mu\neq 0$ can also be solved. Notice that for $\mu=0$, the superconducting chain, in the absence of the coupling to the localized electrons, is in a nontrivial topological state, since the condition $|\mu/2t|<1$ is satisfied [1]. On the other hand, for $\Delta=0$ the model corresponds to the one-dimensional spinless FK model [30].
When $E_{f}=\mu=0$, the Hamiltonian $\mathcal{H_{KL}}=\mathcal{H_{K}}+\mathcal{H_{L}}$ does not depend on the Majorana operators $\beta_{B}$ and $\beta_{A}$, associated with the localized electrons. We can then define an operator $K_{i}=i\beta_{Ai}\beta_{Bi}$ that commutes with the Hamiltonian of the superconductor, i.e., $[K_{i},\mathcal{H_{KL}}]=0$ for all sites. This implies that $K_{i}$ is a $c$-number [26]. On the other hand, the anti-commutation rules of the Majorana operators imply that $K_{i}^{2}=1$ or $K_{i}=\pm 1$. The Hamiltonian of the system can now be written as,
$\mathcal{H}=-\frac{i}{2}\sum_{i}\left[(\Delta+t)\alpha_{Bi}\alpha_{Ai+1}+(\Delta-t)\alpha_{Ai}\alpha_{Bi+1}\right]+i\sum_{i}J_{i}\alpha_{Ai}\alpha_{Bi},$ (8)
where $J_{i}=JK_{i}/16$ ($K_{i}=\pm 1$). From a formal point of view $\mathcal{H}$ is now a quadratic form that can be exactly solved. The interacting problem has been mapped in a non-interacting one, formally in a system with a site dependent chemical potential [35, 36, 38].
Let us consider the case of an infinite chain and introduce the quantities,
@@
where we wrote $J_{i}=Je^{iQr_{i}}$. The homogeneous configuration, with a localized electron on each site of the lattice, $n_{fi}=1$, corresponds to $Q=0$. The staggered configuration, where the values of $K_{i}$ alternate between ±$\pm 1$, such that $n_{fi}$ takes values 00 and $1$ in consecutive sites, corresponds to $Q=\pi$. Notice that $n_{fi}=(1/2)(1+i\beta_{Ai}\beta_{Bi})$. Among all possible configurations, these two play an important role, as we discuss below.
Figure 1: (Color online) Dispersion relations of the excitations in the superconducting state with $t=1$, $\Delta/t=0.2$ and $J/t=0.9$. a) homogeneous case ($Q=0$, $\langle n_{f}\rangle=1$), $\omega_{h}(k)$, and b) staggered case ($Q=\pi$, $\langle n_{f}\rangle=1/2$), $\omega_{s}(k)$. c) energy of the bogoliubons $\omega_{0}(k)$, of the non-interacting case with the same parameters, but $J=0$ (Kitaev chain). The conduction band is always half-filled, and $\mu=E_{f}=0$.
The excitations of the system are given by the k-dependent eigenvalues of the Hamiltonian Eq. 10 [1, 39]. For the homogeneous configuration ($Q=0$), we get,
@@
and $\omega_{h(3,4)}(k)=\pm\omega_{h(1,2)}(k)$. The excitations in the staggered configuration, with $Q=\pi$, are given by
@@
and $\omega_{s(3,4)}(k)=\pm\omega_{s(2,1)}(k)$.
Notice that the condition $\mu=E_{F}=0$ implies that for the non-interacting system, the band of conduction electrons is half-filled. The number of localized $f$-electrons is arbitrary, $n_{f}\leq 1$, and does not affect the total energy of the system, when $J=0$.
In Fig. 1 we plot the dispersion relations of the full problem, Eqs. 11 and 12, for specific values of parameters $\Delta$ and $J$, of the homogeneous and staggered configurations. For completeness, we also show the dispersions of the pure Kitaev chain ($J=0$). In the homogeneous case of Fig. 1a, the gap closes for $J/t=1$, independent of $\Delta$. In the staggered case, Fig. 1b, the gap in the dispersions closes for $J=\Delta$.
In Section IV, to fully describe the topological properties of the interacting system, we investigate numerically the case of finite chains with open boundary conditions. We obtain the spectrum of eigenvalues and the wave functions of the MZMs. However, before that we discuss briefly the pure one-dimensional spinless FK model, since its properties will be relevant for understanding the superconducting case. Notice that this model corresponds to taking the superconducting gap $\Delta=0$ in Eq. 8.
@@ IV The one-dimensional spinless Falicov-Kimball model ($\Delta=0$)
The FK model is a well studied strongly correlated electronic problem, with many applications ranging from heavy fermions to metal-insulator and valence transitions [30]. In particular the one-dimensional spinless FK model has many well-known properties. It can be approached using the technique of bosonisation that yields exact results. From the point of view of topological properties this model is completely trivial.
@@
In this section we briefly discuss the spinless one-dimensional FK model and emphasize some aspects that are relevant when considering the full superconducting problem. Formally, we just take the amplitude of the superconducting pairing $\Delta=0$, to obtain the dispersion relations for the homogeneous and staggered phases of the model. It is possible to distinguish two situations depending on the value of the ratio $(J/t)$.
@@
Figure 2: Dispersion relations of the FK model ($\Delta=0$). Homogeneous case, $\omega_{h}(k)$ for a) $J/t=0.75$, b) $J/t=1$ and c) $J/t=1.2$. Staggered configuration, $\omega_{s}(k)$ for d) $J/t=0.75$, e) $J/t=1$ and f) $J/t=1.2$. In the homogeneous case, a) corresponds to a metallic state and at b) $J/t=1$ there is a metal insulator transition. In the staggered configuration the system is always insulating. The black line is the energy of the localized quasiparticles that coincides with the chemical potential.
For $J/t>1$, both homogeneous and staggered configurations correspond to insulating phases, as shown in Figs. 2c and 2f. Conversely, for $J/t<1$, the homogeneous configuration corresponds to a metallic phase, Fig. 2a, but the staggered one remains insulating, as can be seen in Fig. 2d. Then for the homogeneous case, with one localized electron per site ($n_{f}=1$), there is a metal-insulator transition at $|J/t|=1$, as shown in Fig. 2b for the dispersion relations. The staggered phase is always insulating. Notice that the presence of two dispersive bands is due to an effective hybridization between the localized level and the conduction band due to the FK interaction. If this is too large it can open a gap and give rise to a metal-insulator transition
Due to the condition $E_{f}=\mu=0$, the energy of the pure $f$-electron system does not depend on the number of $f$-electrons, differently from the coupled $s-f$ system. In Fig. 3 we show the ground state energies of the special configurations, $Q=0$ and $Q=\pi$, as functions of the interaction $J$. Notice that the staggered configuration has a lower energy compared to the homogeneous one.
When $\langle n_{f}\rangle=1/2$, besides ordered configurations, as the staggered and other types of superlattices [38], there are many disordered configurations [30, 33], corresponding to a random arrangement of the local moments in the chain. In Fig. 3 we also show the ground state energies, as functions of the interaction $J$, for several realizations of random configurations with $\langle n_{f}\rangle=1/2$. The disordered configurations have a probability $p=1/2$ for a site being occupied, by a localized $f$-electron. The ordered, staggered configuration is always the one with lower energy.
@@
Figure 3: (Color online) Ground state energy of the spinless FK model ($\Delta=0$). Homogeneous case, $n_{f}=1$ (red), staggered, $n_{f}=1/2$ (blue) and random configurations (green) with $\langle n_{f}\rangle=1/2$. Notice that, the staggered configuration has an energy lower than all the probed random configurations with $\langle n_{f}\rangle=1/2$.
@@
V.1 Analysis of the dispersion relations and density of states
We now return to the full problem consisting of a band of spinless fermions, with $p$-wave pairing, that interacts with a background of localized $f$-electrons, Eq. 8. The Hamiltonian with $J=0$, corresponds to a Kitaev chain and an independent collection of localized electrons. On the other hand, for $\Delta=0$, we obtain the FK model.